OpenStudy (anonymous):

Find the y-intercept. 5 + y = -3x y = -3x - 5

OpenStudy (luigi0210):

y-intercept is when x=0

OpenStudy (anonymous):

true

OpenStudy (anonymous):

but i dont get it @Luigi0210

OpenStudy (luigi0210):

Just plug in 0 for x and solve for y.

OpenStudy (anonymous):

i still dont get it

OpenStudy (anonymous):

plug in y into 5+y=-3x

OpenStudy (luigi0210):

It might help if you get your equations in slope-intercept form first.

OpenStudy (anonymous):

so its ganna be 5+0=3(0)